OpenStudy (anonymous):

negative sqaure root 72 *Please show work

OpenStudy (hba):

\[\sqrt{-72}\]

Parth (parthkohli):

Ah, first write this thing as\[\rm i\sqrt{72}\]Now simplify the inner radical.

OpenStudy (hba):

Well @ParthKohli Is Right I Hope You Know How To Do The Rest :)

OpenStudy (hba):

\[i=\sqrt{-1}\]

OpenStudy (anonymous):

i dont

OpenStudy (hba):

Ok I Will SHow You

OpenStudy (hba):



OpenStudy (hba):

\[i \sqrt{2\times2\times2\times3\times3}\]

Parth (parthkohli):

\[\sqrt{-72} = \sqrt{-1} \cdot \sqrt{72}=\rm i\sqrt{72}\]Get it now?

OpenStudy (hba):

\[6i \sqrt{2}\]

OpenStudy (anonymous):

see my thing is i dont know how to solve it or simplify it. i dont know my perfect sqaures eithier.